### v2

# Specific heat of Gd3+- and Eu2+-based magnetic compounds

Abstract
We have studied theoretically the specific heat of a large number of non-frustrated magnetic structures described by the Heisenberg model for systems with total angular momentum $J=7/2$, corresponding to the 4f7 configuration of Gd+3 and Eu+2. For a given critical temperature (determined by the magnitude of the exchange interactions), we find that, to a high degree of accuracy, the specific heat is governed by two primary parameters: the effective number of neighbors, which dictates the extent of quantum fluctuations, and the axial anisotropy. Using these two parameters we fit the specific heat of four Gd compounds and two Eu compounds, achieving a remarkable agreement. Our work opens the possibility of describing the specific heat of 4f7 systems within a general framework.
I Introduction
Rare-earth-based materials are of great interest in both basic and applied condensed matter physics, exhibiting a plethora of captivating fundamental physical phenomena, such as unconventional superconductivity[1, 2], Kondo effect [3, 4, 5], quantum criticality [6], quadrupolar order, and frustration[7, 8, 9, 10, 11].
The specific heat of magnetic materials containing rare earths is typically highly responsive to the environment surrounding the rare earth ions. This sensitivity arises from the splitting of the 4f orbitals by the crystal field and the consequent splitting of the total angular momentum $J$ due to strong spin-orbit coupling. Consequently, excitation energies are influenced by the surrounding environment, thereby affecting the temperature dependence of specific heat.
However, an exception is expected for compounds containing Gd+3 and Eu+2, both of which correspond to the 4f7 configuration. In this case, the ground-state multiplet constructed by Hund’s rules is ${}^{8}S_{7/2}$, which corresponds to total spin $S=7/2$, total orbital angular momentum $L=0$, and total angular momentum $J=7/2$. Since neither the spin nor terms with $L=0$ are affected by crystal fields, one expects that the only source of splitting of the ground-state octuplet is the exchange interaction between rare-earth ions responsible for the magnetic structure of the compound. This is true as a first approximation. In fact, we will show that the specific heat of five Gd compounds can be well fitted ignoring crystal-field effects.
Nevertheless, due to spin-orbit coupling, the ground-state multiplet acquires a small admixture of the ${}^{6}P_{7/2}$ state with $L=1$. An explicit calculation for Gd+3 presented in Appendix B of Ref. 12 estimates that the excited multiplet contributes approximately $3\%$ to the ground state. This admixture introduces a component with $L=1$ in the ground state, allowing the ground state to split under the influence of crystal fields if the symmetry at the rare-earth site is not high, with only a single axis of higher order. In such cases, the anisotropy term $H_{A}=C[3L^{2}_{z}-L(L+1)]$ becomes relevant. Note that in high-symmetry environments such as tetragonal, cubic, or octahedral point groups ($T$, $T_{d}$, $T_{h}$, $O$, $O_{h}$), the $L=1$ states remain unsplit. This result, derived from group theory, is intuitively understood by noting that a basis for $L=1$ states can be chosen to transform like the unit vectors $\hat{x}$, $\hat{y}$, $\hat{z}$, which are equivalent in high-symmetric environments. In contrast, if the maximum order axis of the point group is 2, a term $L^{2}_{x}-L^{2}_{y}$ is allowed. In even lower symmetries, $L_{x}L_{y}$ may also arise; however, their intensity is expected to be much smaller, and we can fit experimental data without including them.
Using the Wigner-Eckart theorem, $H_{A}$ can be written in the form $H_{A}=K[3J^{2}_{z}-J(J+1)]$, where $K$ is of order $\sim 3\%C$.
Even if $H_{A}$ can be neglected, systematic deviations of the observed magnetic contribution to the specific heat $C(T)$ in Gd-based compounds from the corresponding mean-field result have been observed since long time ago, pointing to the importance of quantum fluctuations. Early measurements of $C(T)$ on GdNi5 [13] were compared with calculations in the molecular field approximation for $T<T_{C}=32$ K, where $T_{C}$ is the critical temperature, with good agreement. However, the authors recognized evident discrepancies at low temperature and in the vicinity of $T_{C}$ that can be attributed to “short-range ordering effects not taken into account”. Simultaneous measurements of $C(T)$ on other Gd compounds [14] show more significant deviations from the mean-field result. The origin of such deviations was investigated by introducing modulations in the amplitude of the magnetic moments in the mean-field approximation [15]. Such a model describes different possibilities of the observed temperature dependence of $C(T)$ for $T<T_{C}$. However, some compounds clearly escape from those predictions, especially concerning observed magnetic fluctuations right above $T_{C}$.
Recent research on Eu2+-based compounds shows similar shortcomings of the mean-field approach [16, 17]. The fact that Eu2+ has a much larger atomic volume and therefore may participate in the formation of very different crystalline structures indicates that the observed deviations are an intrinsic characteristic of the magnetic interactions and not dependent on crystalline symmetries or the sign of the interactions. Therefore, for a quantitative explanation of the observed $C(T)$, it is necessary to include quantum fluctuations and go beyond mean field. The effect of these fluctuations was calculated recently using cluster mean-field theory [18].
In this work, we report the specific heat of a large number of structures described by a Heisenberg Hamiltonian $H$ containing exchange interactions at different distances and an axial anisotropy term $H_{A}$:
$H=H_{0}+H_{A},$ (1)
where
$H_{0}=\sum_{i,\delta}J_{\delta}\mathbf{S}_{i}\cdot\mathbf{S}_{i+\delta}/2,\text{ and }H_{A}=K\sum_{i}[3S_{iz}^{2}-S(S+1)].$ (2)
$\mathbf{S}_{i},S_{iz}$ denote the spin (or angular momentum) $|S|=7/2$ operators at site $i$ and its projection, $\delta$ labels the non-equivalent space vectors connecting different spins at short distances (nearest and possibly further neighbors), and $J_{\delta}$ is the corresponding exchange interaction. We use $S$ instead of $J$ to denote the operators to avoid confusion with the exchange interactions.
The theoretically studied magnetic structures are non-frustrated and correspond to either ferromagnetic interactions or non-frustrated antiferromagnetic ones. We find that the specific heat of all these structures can be characterized to a large degree of accuracy by two parameters: i) the effective number of neighbors (to be defined more precisely below) and ii) the magnetic anisotropy defined by $H_{A}$. Using these two parameters, we fit the reported specific heat as a function of temperature for six compounds, obtaining an impressive agreement.
The effect of fluctuations on the specific heat of a spin system described by the Heisenberg model is expected to decrease as the number of neighboring spins increases. We can draw an analogy to a binary random distribution (which would correspond to spin 1/2). It is known that after $N$ attempts in a binary random distribution $x$ with probabilities $p$ and $1-p$ for the values $x=1$ and $x=-1$, respectively, the mean value (proportional to the magnetic moment) is $\langle x\rangle=(2p-1)N$, and the standard deviation is $\sqrt{\langle x^{2}\rangle-\langle x\rangle^{2}}=2\sqrt{Np(1-p)}$. Therefore, the ratio between the standard deviation and the mean value decreases as $1/\sqrt{N}$. Similarly, in our quantum Heisenberg model, one expects that as the number of neighbors increases, the specific heat is more similar to the mean-field result, in which quantum fluctuations are neglected and the effective magnetic field that each ion feels is replaced by its mean value.
Generalizing the above results to the case of a distribution with intensities ±$\pm|J_{\delta}|$ for $z_{\delta}$ neighbors, we obtain after some algebra (see Appendix):
$\frac{\langle x^{2}\rangle-\langle x\rangle^{2}}{\langle x\rangle^{2}}=\frac{4p(1-p)}{(2p-1)^{2}}r,\quad\text{where}\quad r=\frac{\sum_{\delta}z_{\delta}J_{\delta}^{2}}{\left(\sum_{\delta}z_{\delta}|J_{\delta}|\right)^{2}}=\frac{1}{z},$ (3)
Here, $z_{\delta}$ represents the number of neighbors of type $\delta$ (with exchange interaction $J_{\delta}$), and the summation goes over all types of neighbors. The quantity $z=1/r$ captures the effective number of neighbors, taking into account the strength ($J_{\delta}^{2}$) and number ($z_{\delta}$) of each type of interaction. Remarkably, for unfrustrated systems, the observed specific heat depends mainly on this effective number of neighbors $z$, independent of the specific lattice structure. For larger $z$ (higher coordination number), the results approach the mean-field ones, where quantum fluctuations are suppressed. Conversely, for lower $z$ (fewer neighbors), quantum fluctuations become more important, leading to a smaller specific heat below the critical temperature $T_{C}$ and a larger specific heat above $T_{C}$.
Moreover, for some actual systems, the effect of the axial anisotropy $H_{A}$ becomes important, redistributing the entropy below $T_{C}$ (see Section IV).
Using the above two parameters for simple lattices and the known value of $T_{C}$, we have fitted the data for the specific heat of GdNi3Ga9, GdPbBi, GdCu2Ge2, GdNiSi3, Eu2Pd3Sn3, and EuPdSn2. The corresponding results are presented in Section VI. Using essentially the same method, the specific heat of GdCoIn5 has been fitted in Ref. 19.
II Methods
We have calculated the specific heat of a large number of magnetic structures from the numerical derivative of the internal energy as a function of temperature. The internal energy was calculated using quantum Monte Carlo (QMC) simulations with the “minbounce” algorithm (“dirloop_sse” application) from the ALPS libraries [20, 21]. Simulations were performed on systems of up to $8\times 30^{3}$ magnetic moments.
To generate the lattice configurations for the ALPS simulations, we utilized the “SUNNY” package [22] in conjunction with crystallographic information files (CIF) obtained from public repositories. Specifically, CIF files for Fe in simple cubic (SC), body-centered cubic (BCC), hexagonal close-packed (HCP), and face-centered cubic (FCC) structures, as well as C in diamond, and MgAl2Cu (for EuPdSn2) were retrieved from the Materials Project [23] database. The ErNi3Al9 CIF file was obtained from the Crystallography Open Database [24] using data from reference 25. In each case, the magnetic structure was created by replacing C, Fe, Mg, or Er with a spin moment $S=7/2$.
For all lattices, we adopted a rectangular cuboid supercell geometry as the conventional unit cell, resulting in varying numbers of sites within the supercell, with the diamond lattice requiring the largest (8 sites). Simulations were performed on systems containing $N^{3}$ supercells with periodic boundary conditions, typically with $N=10$, $20$, and $30$. For the critical temperature, results were extrapolated to the thermodynamic limit assuming a $1/N$ dependence.
The well-known sign problem [26] associated with QMC simulations limited our investigation to non-frustrated magnetic structures. At low temperatures, the number of Monte Carlo steps required for statistically reliable results increases exponentially. Consequently, the internal energy was primarily computed in a temperature range between $0.2T_{C}$ and $2T_{C}$, where $T_{C}$ is the critical temperature of the system. In a few cases, simulations were extended to $0.1T_{C}$ at low temperatures. However, the reliability of the results below $0.2T_{C}$ is limited. We used $2\times 10^{5}$ Monte Carlo steps for thermal equilibration and $6\times 10^{5}$ steps for measurements.
The coordination number (number of neighbors included) for each lattice is provided in parentheses after its name. For example, FCC(18) refers to a face-centered cubic lattice with 12 first neighbors and 6 second-nearest neighbors. Only for the simple cubic case, SC(14), we consider a scenario including both the 8 nearest neighbors (NN) and the 6 third NN. In other cases, additional neighbor shells were included by increasing the considered neighbor radius $d$.
Our theoretical results for the specific heat for all lattices used here are available in 27.
III Role of the effective number of neighbors
Figure 1: Critical temperature as a function of the number of neighbors for several ferromagnetic structures, sizes ($N=10,20,30$) and extrapolated value. The straight line denotes the mean-field value $T_{C}^{\text{MF}}$.
In this Section we report our results for the specific heat and critical temperature for several simple lattices as a function of the number of neighbors that interact with a given site. For simplicity, in all lattices considered, all sites are equivalent, and we assume all $J_{\delta}=J_{1}$, where $J_{1}$ is the NN interaction, up to a certain distance $d$ and 0 for longer distances. Therefore, the effective number of neighbors $z=\Sigma_{\delta}z_{\delta}$ is just the sum of all neighbors within a sphere of radius $d$. In Section V we discuss more general cases showing that the results are not severely affected by this assumption.
Most of our studies assume ferromagnetic interactions ($J_{\delta}<0$). Generally, the results for the corresponding non-frustrated antiferromagnetic structures are very similar, particularly for $z>6$, except at very low temperatures, where the specific heat dominated by magnons behaves as $T^{3/2}$ ($T^{3}$) for ferro- and antiferromagnetic structures, respectively.
In Fig. 1, we show the critical temperature $T_{C}$ as a function of $z$ for the different structures considered. Calculations have been done in finite systems containing $N^{3}$ conventional unit cells, with $N=10$, 20, and 30, as explained in Section II. The extrapolated value is also indicated in the figure.
As expected, $T_{C}$ increases almost linearly with $z$ as the mean-field value $T_{C}^{\text{MF}}$ (the effective magnetic field due to the interactions is proportional to $zJ_{1}$, see Section IV.1), but it is lower than $T_{C}^{\text{MF}}$, particularly for low $z$ due to the effect of quantum fluctuations, which is larger for low $z$.
For large enough $z$ one expects that $T_{C}$ and $T_{C}^{\text{MF}}$ coincide. However, for large $z$ in a finite system with periodic boundary conditions, the effective coordination is smaller when the range of the interactions $d$ becomes of the order of the system size, because neighbors at different distances coincide, and it is not possible to reach this limit with our method.
Figure 2: (Color online) Specific heat over temperature as a function of temperature for several ferromagnetic structures with different effective number of neighbors $z$. Top: $z\leq 12$. Bottom: $z\geq 12$. There is an almost perfect superposition between HCP and FCC for $z=12$, among SC, FCC and HCP for $z=18$, and among SC, BCC and the EuPdSn2 structure for $z=26$. Therefore, only one of the curves for each of these three sets is shown. LHc(4) corresponds to the layered honeycomb structure with $z=4$. Bottom inset shows the specific heat for FCC(12) lattices with sizes $N=10$ (black line), $20$ (red), and $30$ (green) in the range $1.50.5<T/T_{C}<$ 1.5.
In Fig. 2 we show the results for the specific heat as a function of temperature for a large number of ferromagnetic structures with different numbers of magnetic neighbors $z$. Also shown is the mean-field result which corresponds to the limit $z\rightarrow\infty$. This result shows larger $C$ for $T<T_{C}$, while $C=0$ for $T>T_{C}$. In contrast, the specific heat for a honeycomb layered system with $z=4$ (three nearest neighbors within the plane of the honeycomb layer and one above or below the layer in alternating sites) shows the smallest specific heat below $T_{C}$ and the largest one above $T_{C}$. Note that the diamond structure, which is completely different from the previous one but also has $z=4$, has a very similar specific heat. The low-temperature shoulder is a Schottky-like effect arising from the internal field [19, 15]. The lower panel inset shows that finite size effects are only appreciable close to $T_{C}$ as exemplified in the case of the FCC(12) lattice. For this reason, we show specific heat curves from QMC corresponding to $N=20$.
Increasing $z$ from 4 to 6, the specific heat of the simple cubic structure with six NN displays the expected trend of increasing (decreasing) $C$ below (above) $T_{C}$. Increasing $z$ up to 14 using the BCC structure with 6 NN and 8 next-nearest neighbors (NNN), the same trend is observed in general. However, the specific heat of the honeycomb lattice with $z=10$ (four links as described before and adding 6 second neighbors in the plane) lies slightly below the cubic BCC with $z=8$ nearest neighbors for $T\sim 0.5T_{C}$. In addition, somewhat unexpectedly, the result for the simple cubic structure including 6 NN and 8 third NN, shows that $C$ at $T\sim 0.5T_{C}$ is greater than the corresponding one of the above-mentioned BCC structure with the same $z=14$, and also greater than the specific heat of one structure with $z=16$ and two structures with $z=18$. Nevertheless, except for this case and the above-mentioned inversion between $z=8$ and 10, the specific heat is very similar for different structures with the same $z$ and increases with increasing $z$ at intermediate temperature $T\sim 0.5T_{C}$.
IV Effect of anisotropy
In this Section, we address the role of anisotropy first in the mean-field approximation and then incorporate the effect of quantum fluctuations.
IV.1 Mean-field Approximation
In this subsection, we discuss the effect of the term $H_{A}$ on the specific heat calculated in the mean-field approximation, in which it is assumed that each site is subjected to a constant effective magnetic field resulting from the interaction with the remaining sites. Specifically, the interactions appearing in $H_{0}$ [see Eq. (1)] are approximated as
$\mathbf{S}_{i}\cdot\mathbf{S}_{j}\approx\langle\mathbf{S}_{i}\rangle\cdot\mathbf{S}_{j}+\mathbf{S}_{i}\cdot\langle\mathbf{S}_{j}\rangle-\langle\mathbf{S}_{i}\rangle\cdot\langle\mathbf{S}_{j}\rangle,$ (4)
We assume that all sites are equivalent and there is no external magnetic field. For simplicity, we restrict the analysis to the easy axis case $K<0$ for which the spins point in the direction of the anisotropy axis $z$. In the next subsection, both signs of $K$ and the effects of quantum fluctuations will be considered. For an antiferromagnetic order, we rotate 180 degrees the spins pointing in the $-z$ direction around a perpendicular axis so that for all sites $\langle S_{z}\rangle\geq 0$. This expectation value serves as the order parameter for the magnetic phase.
The effective magnetic field $B_{\text{eff}}$ is given by
$\tilde{B}=g\mu_{B}B_{\text{eff}}=I\langle S_{z}\rangle\text{,~{} }\langle S_{z}\rangle=\sum\limits_{M=-J}^{J}\frac{Me^{-\beta E_{M}}}{Z}\text{,~{}}Z=\sum\limits_{M=-J}^{J}e^{-\beta E_{M}},$ (5)
where $g$ is the gyromagnetic factor, $\mu_{B}$ is the Bohr magneton, $I=|\Sigma_{\delta}z_{\delta}J_{\delta}|$ is the sum of all interactions ($z_{\delta}$ is the number of neighbors at distance $\delta$ and $J_{\delta}$ its intensity), $\beta=1/(k_{B}T)$ where $T$ is the temperature and $k_{B}$ is the Boltzmann constant. The energies $E_{M}$ for each eigenvalue $M$ of the operator $S_{z}$ are given by:
$E_{M}=-\tilde{B}M+K[3M^{2}-S(S+1)].$ (6)
Eqs. (5) and (6) permit to determine $\tilde{B}$ self-consistently. In the limit $\tilde{B}\rightarrow 0$ one obtains an equation for the critical temperature
$k_{B}T_{C}^{\text{MF}}=I\langle S^{2}_{z}\rangle\text{, where~{}~{}}\langle S^{2}_{z}\rangle=\sum\limits_{M=-J}^{J}\frac{M^{2}e^{-\beta E_{M}}}{Z}.$ (7)
The specific heat is obtained from the numerical derivative of the entropy $S$ with respect to temperature:
$C=T\frac{dS}{dT}\text{,~{}~{}~{} }S=\sum\limits_{M=-J}^{J}p_{M}\ln p_{M}\text{,~{}~{}~{} }p_{M}=\frac{e^{-\beta E_{M}}}{Z}$ (8)
Figure 3: (Color online) Specific heat over temperature as a function of temperature for several anisotropies in the mean-field approximation.
In Fig. 3 we show the resulting effect of the axial anisotropy on the specific heat keeping the value of $k_{B}T_{C}=1$ constant, as the unit of energy. For $K=0$, the value of the intensity for a given $T_{C}$ is $I=3k_{B}T_{C}/[S(S+1)]$, leading to $I=4/21\sim 0.190$ for $k_{B}T_{C}=1$.
For $K=-0.02$, we obtain $I=0.1531$ from Eq. (7). This is because the expectation value $\langle S^{2}_{z}\rangle$ that enters Eq. (7) is larger for $K<0$ and should be compensated with a smaller $I$. In addition, since the values of high $S_{z}$ projection $|M|$ are favored, the effective magnetic field is more efficient than for $K=0$ in splitting the lowest energy levels and therefore the entropy and specific heat decrease for $T<0.4T_{C}$ with respect to the isotropic case. This is compensated by an increase in $C/T$ for $0.5T_{C}<T<T_{C}$.
For $K=-0.04$, $I$ decreases further to $I=0.1263$, and the changes in the specific heat are more marked in the same direction. For large negative $K$, $I=1/(7/2)^{2}\sim 0.0816$.
IV.2 Beyond Mean-field
In this subsection we discuss the effect of anisotropy in cases where the fluctuations of the effective field are also included.
Figure 4: (Color online) Specific heat over temperature as a function of temperature for several anisotropies for the simple cubic structure including 6 NN and 8 third NN with $J_{\delta}=-1$.
In Fig. 4 we show how the specific heat for a structure with $z=14$ is modified by the addition of $H_{A}$. The results are qualitatively the same as in the mean-field case for $T<T_{C}$. Addition of the anisotropy term shifts the entropy from the region $T<0.5T_{C}$ mainly to the region $0.5T_{C}<T<T_{C}$, and the effect is more pronounced for negative $K$.
However, in contrast to the mean-field case, addition of the anisotropy term decreases the specific heat for $T>T_{C}$. This reduction is expected because the specific heat in this region primarily results from quantum fluctuations, which are diminished by the energy level splitting induced by $H_{A}$.
V Comparison of different structures with the same effective number of neighbors
Figure 5: (Color online) Specific heat over temperature as a function of temperature for several structures with different values of $r$ defined by Eq. (3). Top left: several structures with $J_{\delta}=-1$ and different $r$. The rest of the panels show the effect of changing $J_{\delta}$ keeping the same $r$.
In Section III, we discussed the effect of increasing the number of neighbors $z$ in the specific heat, while keeping $J_{\delta}=J_{1}$. Here we investigate the effect of changing $J_{\delta}$ for first, second, and possibly third nearest neighbors, keeping the same $r$ [see Eq. (3)], which corresponds to the inverse of the effective number of neighbors $z$. Some parameter sets may be highly unrealistic. However, the objective of this section is to demonstrate that entirely different parameter sets, which result in the same effective $z$, produce remarkably similar specific heat behaviors.
In Fig. 5 (b) we show the comparison for a moderately small $z$: a cubic BCC structure with 6 NN and 8 third NN. We compare the case of NN interaction $J_{1}=-1$ and NNN interaction $J_{2}=0$ with $J_{1}=-1/8$ and $J_{2}=-1$. In spite of the radical change in the magnitude of the interactions, the specific heats are similar.
In Fig. 5 (c), we show a similar comparison for an FCC structure with different combinations of $J_{1}$ and $J_{2}$. Again the results are similar, except for the case in which the dominant interaction is changed drastically from NN to NNN. When the dominant interaction is that of NNN, the specific heat is larger and the difference is of the order of that between different $z$ for $J_{1}=-1$ (Fig. 5 (a)). Finally, in Fig. 5 (d), we show a similar comparison for a larger $z$. In this case, the agreement between the two curves is excellent.
In conclusion, except for extreme cases, the magnitude of the specific heat at $T\sim T_{C}/2$ is mainly determined by the effective number of neighbors $z=1/r$. This result is significant, because it allows us to predict the expected specific heat taking into account the effective number of neighbors with a relative independence of the detailed magnetic structure.
VI Fits of experimental data
In this Section, we present fits of the specific heat of six compounds: GdNi3Ga9, GdPbBi, GdCu2Ge2, GdNiSi3, Eu2Pd3Sn3, and EuPdSn2. Previously, the specific heat of GdCoIn5 has been fitted [19].
The fitting parameters are the known $T_{C}$, which determines the magnitude of the interactions $J_{\delta}$ assumed equal, the number of neighbors $z$ and for the Eu compounds, the anisotropy $K$. The Gd compounds could be fitted with $K=0$. The assumption of $J_{\delta}$ independent of $\delta$ is justified from the results of the previous Section, which show that different $J_{\delta}$ with the same effective number of neighbors $1/r$ show very similar specific heats. The fact that $K$ is important only for the Eu compounds, might be expected for the $\sim 25\%$ larger ionic radius of Eu2+ compared to that of Gd3+ [28], rendering the former more sensitive to crystal-field effects.
The results are presented in Fig. 6. The Gd compounds are ordered with increasing effective number of neighbors $z$ in the fit. The parameters of the fits used for the different compounds are listed in Table 1. Also indicated is the number of NN in the reported structure, denoted as $z_{0}$. Notations like 1+1+2 in $z_{0}$ indicate that there are four atoms located at three nearly identical distances.
Compound $T_{N}$ (exp,[K]) $z_{0}$ $z$(fit) $J$[K] $K$[K]
GdNi3Ga9 20 3 4 1.52 0
GdCu2Ge2 12 4 8 0.38 0
GdPbBi 13 12 10 0.33 0
GdNiSi3 22 2+2+2 26 0.18 0
Eu2Pd3Sn 47 1+1+2 14 0.66 -0.26
EuPdSn2 12 2+2 14 0.55 -0.44
Table 1: Parameters used in the fits: $T_{N}$ is the Neel temperature, $z_{0}$ is the number of nearest neighbors in the reported structure, $z$(fit) is the effective number of neighbors in the fit, $J$ is the intensity of the interactions, and $K$ is the anisotropy.
GdNi3Ga9 [29, 30] clearly corresponds to a low coordination number ($z=4$). This result is expected, since the honeycomb lattice described in Section III with antiferromagnetic interactions (three intralayer and one interlayer for each site) corresponds to the structure of the material and provides a very good fit.
For the rest of the structures, ferromagnetic interactions are assumed. As stated in Section III, as long as there is no significant frustration, the difference between ferromagnetic and antiferromagnetic interactions for $z>6$ is only noticeable at very low temperatures, well below the temperature range of interest in this study. The key features of the specific heat curves we analyze, such as the shape of the peak and the high-temperature tail, are primarily determined by the effective coordination number and the single-ion anisotropy, and these features are largely insensitive to the sign of the exchange interaction in the absence of strong frustration.
GdCu2Ge2 has a tetragonal structure near to a BCC one (Fig. 4.1 of Ref. [31]). Assuming superexchange interaction through Cu, the “BCC” like character is even more plausible. Using only 8 NN interactions of a BCC structure, the resulting specific heat is very near the experimental one, although slightly below (above) the experimental one for $T<T_{C}$ ($T>T_{C}$), suggesting that $z$ might be a little bit larger. If only the four NN were considered, the system would be two-dimensional (with a broadened transition near $T_{C}$) and this is incompatible with the observed specific heat.
GdPdBi [32] is a half-Heusler structure with antiferromagnetic coupling. We obtain a very good fit assuming a layered honeycomb lattice as for GdNi3Ga9, but with a larger number of effective neighbors (10). The fitting is better than using an FCC structure. This effect might be caused by a small distortion of the lattice. Note that in this case, the effective number of neighbors ($z=10$) is slightly lower than the number of nearest neighbors ($z_{0}=12$). This is likely due to a slight frustration, which enhances the relative impact of fluctuations.
For GdNiSi3 [33], a simple cubic structure with $z=26$ including neighbors up to the third coordination shell provides a very good fit of the data. For a three-dimensional behavior, $z>16$ is necessary to connect different planes.
This completes the Gd compounds that we have considered. In a previous work [19], the specific heat of GdCoIn5 has been fitted with essentially the same procedure. The compound has a tetragonal structure. The authors have extrapolated to the thermodynamic limit, the specific heat that corresponds to a simple cubic lattice $z=6$ NN with antiferromagnetic interactions, providing an excellent fit of the data.
Now turning to Eu compounds, the specific heat of Eu2Pd3Sn [34] is very well fitted using a simple cubic structure with 6 NN and 8 third NN ($z=14$) with $J_{1}=J_{2}$ and $K=-0.4J_{1}$. Small deviations exist at very low temperatures, where our numerical method is not fully reliable. The point group at the Eu sites is $C_{1h}$, compatible with the presence of $H_{A}$ as discussed in the introduction.
Finally, the same simple cubic structure, but with $K=-0.8J_{1}$ provides a very good fit of the specific heat of EuPdSn2 [35]. For temperatures very near $T_{C}$, the theoretical data are a little above (below) the experimental ones for $T<T_{C}$ ($T>T_{C}$). The point group at the Eu sites is $C_{2v}$, which also allows for the inclusion of the term $H_{A}$. The discrepancies near $T_{C}$ ($T_{N}$) likely originate from factors not explicitly included in our model, such as more complex magnetic ordering (e.g., a collinear amplitude-modulated structure) resulting from modulated interactions and crystal-field effects, as discussed in Ref. 15.
Since all the compounds mentioned above contain transition-metal ions, which in some cases are magnetic, one might question whether they influence the reported specific heat. However, our analysis shows that the entropy obtained by integrating $C/T$ is very near ln(8) per rare-earth ion (consistent with $S=7/2$). This indicates that any effect from the transition metals is negligible.
Figure 6: (Color online) Specific heat over temperature as a function of temperature for several compounds and our fit. $K=0$ for the Gd compounds. Experimental data were taken from Refs. 29, 30, 32, 31, 33, 35, 34.
VII Summary
Using quantum Monte Carlo, we have calculated the specific heat of a large number of non-frustrated magnetic structures. As expected from mean-field theory, the critical temperature $T_{C}$ increases approximately proportionally to $I=\Sigma_{\delta}z_{\delta}|J_{\delta}|$, where $J_{\delta}$ is the exchange interaction at distance $\delta$ and $z_{\delta}$ is the number of neighbors at this distance. However, $T_{C}$ lies below the mean-field value.
Rather surprisingly, we find that the specific heat depends mainly on the effective number of neighbors $z$ [see Eq. (3)] and is not quite sensitive to the particular structure. This allows us to derive general conclusions on the effect of $z$, independently of the particular structure. For large $z$, quantum fluctuations are reduced, and the specific heat approaches the mean-field value. Reducing $z$, the specific heat decreases below $T_{C}$ and increases above $T_{C}$ displaying larger tails.
For a quantitative comparison with experimental results, it is essential to incorporate axial anisotropy $H_{A}$ (see Section IV). The origin of $H_{A}$ lies in a small admixture of ${}^{6}P_{7/2}$ in the ground-state multiplet. This admixture introduces a component with $L=1$, rendering the system sensitive to axial crystal fields.
Using $z$ as the only free parameter in addition to the critical temperature, we have fitted the specific heat of four Gd compounds (GdNi3Ga9, GdPdBi, GdCu2Ge2, and GdNiSi3). The specific heat of GdCoIn5 has been fitted previously using the same approach [19]. Curiously, we find that $H_{A}$ can be neglected for these Gd compounds. Including also $K$ as a parameter, we have done the same fitting for two Eu compounds (Eu2Pd3Sn3 and EuPdSn2). The agreement shown in Fig. 6 is noticeable. It is possible that the larger atomic volume of Eu+2 renders this ion more susceptible to crystal-field effects.
Note that while most of our calculations have been done on non-frustrated systems, some of the fitted compounds here (GdPdBi and GdNiSi3), and GdCoIn5 fitted previously [19], are expected to have some degree of frustration.
Our work opens the possibility to classify and analyze the specific heat of non-frustrated Gd+3 and Eu+2, at least in a first semiquantitative manner using the above-discussed two parameters.
Acknowledgements.
The authors are grateful to I. Čurlík, L. de Sousa Silva, J.G.S. Duque, P.G. Pagliuso, and M. Avila for allowing us access to the original experimental data. A.A.A. acknowledges financial support provided by PICT 2020A 03661 of the Agencia I+D+i, Argentina. D.J.G. is supported by PIP 11220200101796CO of CONICET, Argentina.
VIII Appendix
Here we calculate the mean value and fluctuations of a random variable $x$ with different intensities  [^1] The same result can be obtained considering each bond as an independent variable.. For simplicity we consider only two intensities $J_{1}$ and $J_{2}$. Extension to the general case is straightforward. We consider $z_{1}$ attempts in which the probability of the result $J_{1}$ is $p$ and that of $-J_{1}$ is $q=1-p$. In addition, there are $z_{2}$ attempts with probability $r$ of finding the result $J_{2}$ and $s=1-r$ for the result $-J_{2}$. At the end we will take $r=p$. For non-frustrated systems, both $J_{\delta}$ can be chosen positive. Clearly, the expectation value of $x$ is
$\langle x\rangle=z_{1}J_{1}(p-q)+z_{2}J_{2}(r-s).$ (9)
On the other hand, considering the probability of each individual event, we have:
$\langle x^{2}\rangle$ $=$ $\sum\limits_{n=0}^{z_{1}}\sum\limits_{m=0}^{z_{2}}\binom{z_{1}}{n}\binom{z_{2}}{m}p^{n}q^{z_{1}-n}r^{m}s^{z_{2}-m}$ (10)
×$\times\left[nJ_{1}-(z_{1}-n)J_{1}+mJ_{2}-(z_{2}-m)J_{2}\right]^{2}.$
Using the operator
$O=\left[J_{1}\left(p\frac{\partial}{\partial p}-q\frac{\partial}{\partial q}\right)+J_{2}\left(r\frac{\partial}{\partial r}-s\frac{\partial}{\partial s}\right)\right],$
[…]
and a similar expression for $(r+s)^{z_{2}}$, where $p$ and $q$ as well as $r$ and $s$ are taken as independent variables in the derivatives, one realizes that the expectation value Eq. (10) can be written as
$\langle x^{2}\rangle=O^{2}(p+q)^{z_{1}}(r+s)^{z_{2}}.$ (12)
Performing the calculation, we find after some algebra
$\langle x^{2}\rangle$ $=$ $J_{1}^{2}z_{1}\left[1+(z_{1}-1)(p-q)^{2}\right]+J_{2}^{2}z_{2}\left[1+(z_{2}-1)(r-s)^{2}\right]$ (13)
$+2J_{1}J_{2}z_{1}z_{2}(p-q)(r-s)$
Using Eqs. (9) and (13), setting $r=p$ and replacing $q=s=1-p$, one finally obtains
$\frac{\langle x^{2}\rangle-\langle x\rangle^{2}}{\langle x\rangle^{2}}=\frac{4p(1-p)}{(2p-1)^{2}}\frac{z_{1}J_{1}^{2}+z_{2}J_{2}^{2}}{\left(z_{1}J_{1}+z_{2}J_{2}\right)^{2}}.$ (14)
Naturally, if $\langle x\rangle=0$ as it is the case for temperatures above the critical temperature $T_{C}$, Eq. (14) gives a divergent result. In this case a meaningful comparison should involve the fluctuations above $T_{C}$ and the value of $\langle x\rangle=0$ at some point below $T_{C}$.