### v3

# Specific heat of Gd3+ and Eu2+-based magnetic compounds

Abstract
We have studied theoretically the specific heat of a large number of non-frustrated magnetic structures described by the Heisenberg model for systems with total angular momentum $J=7/2$, corresponding to the 4f7 configuration of Gd+3 and Eu+2. For a given critical temperature (determined by the magnitude of the exchange interactions), we find that, to a high degree of accuracy, the specific heat is governed by two primary parameters: the effective number of neighbors, which dictates the extent of spatial and quantum fluctuations, and the axial anisotropy. Using these two parameters we fit the specific heat of four Gd compounds and two Eu compounds, achieving a remarkable agreement. Our work provides a framework to describe the specific heat of 4f7 systems within a unified framework.
I Introduction
Rare-earth-based materials are of great interest in both basic and applied condensed matter physics. These materials exhibit a wide range of fascinating physical phenomena. For example, they show unconventional superconductivity[1, 2], Kondo effect [3, 4, 5], quantum criticality [6], quadrupolar order and frustration[7, 8, 9, 10, 11].
The specific heat of magnetic materials containing rare earths is typically highly sensitive to the environment surrounding the rare earth ions. This sensitivity arises from the splitting of the 4f orbitals by the crystal field and the consequent splitting of the total angular momentum $J$ due to strong spin-orbit coupling. Consequently, excitation energies are influenced by the surrounding environment. This, in turn, affects the temperature dependence of specific heat.
However, an exception is expected for compounds containing Gd+3 and Eu+2. Both of these ions correspond to the 4f7 configuration. In this case, the ground-state multiplet, constructed by Hund’s rules, is $^{8}S_{7/2}$. This corresponds to a total spin $S=7/2$, total orbital angular momentum $L=0$, and total angular momentum $J=7/2$. We show in this manuscript that the specific heat of 4f7 systems can be characterized primarily by two parameters: the effective number of neighbors $z$ and axial anisotropy $K$. This effective number of neighbors $z$ accounts for the number of interaction bonds of each ion. This simple framework could enable more accurate predictions of thermal properties in magnetic materials and serve as guide for specific model Hamiltonians.
Early measurements of $C(T)$ on GdNi5 [12] were compared with calculations in the molecular field approximation for $T<T_{C}=32$ K, where $T_{C}$ is the critical temperature, with good agreement. However, the authors recognized evident discrepancies at low temperature and in the vicinity of $T_{C}$. These discrepancies were attributed to “short-range ordering effects not taken into account”. Simultaneous measurements of $C(T)$ on other Gd compounds [13] show more significant deviations from the mean-field result. The origin of such deviations was investigated by introducing modulations in the amplitude of the magnetic moments within the mean-field approximation [14]. Such a model describes different possibilities of the observed temperature dependence of $C(T)$ for $T<T_{C}$. However, some compounds clearly escape from those predictions. This is especially true for observed magnetic fluctuations right above $T_{C}$.
Recent research on Eu2+-based compounds shows similar shortcomings of the mean-field approach [15, 16]. The fact that Eu2+ has a much larger atomic volume suggests it may participate in the formation of distinct crystalline structures. This indicates that the observed deviations are an intrinsic characteristic of the magnetic interactions. They are not dependent on crystalline symmetries or the sign of the interactions. Therefore, for a quantitative explanation of the observed $C(T)$, it is necessary to include spatial and quantum fluctuations and go beyond mean field. The effect of these fluctuations was calculated recently using cluster mean-field theory [17].
In this work, we report the specific heat of a large number of structures described by a Heisenberg Hamiltonian $H$ containing exchange interactions at different distances and an axial anisotropy term $H_{A}$:
$H=H_{0}+H_{A},$ (1)
where
$H_{0}=\sum_{i,\delta}J_{\delta}\mathbf{S}_{i}\cdot\mathbf{S}_{i+\delta}/2$ (2)
and
$H_{A}=K\sum_{i}[3S_{iz}^{2}-S(S+1)].$ (3)
$\mathbf{S}_{i},S_{iz}$ denote the spin (or angular momentum) $|S|=7/2$ operators at site $i$ and its projection, $\delta$ labels the non-equivalent space vectors connecting different spins at short distances (nearest and possibly further neighbors), and $J_{\delta}$ is the corresponding exchange interaction. Couplings $J_{\delta}$ are ferromagnetic if negative or antiferromagnetic if positive. We use $S$ instead of $J$ to denote the operators to avoid confusion with the exchange interactions.
The theoretically studied magnetic structures are non-frustrated and correspond to either ferromagnetic interactions or non-frustrated antiferromagnetic ones. We find that the specific heat of all these structures can be characterized to a large degree of accuracy by two parameters: i) the effective number of neighbors (to be defined more precisely below) and ii) the magnetic anisotropy defined by $H_{A}$. Using these two parameters, and the known value of $T_{C}$, we fit the data for the specific heat of GdNi3Ga9, GdPbBi, GdCu2Ge2, GdNiSi3, Eu2Pd3Sn3, and EuPdSn2. The corresponding results are presented in Section VI.
I.1 The Effective Number of Neighbors
The effect of fluctuations on the specific heat of a spin system described by the Heisenberg model is expected to diminish as the number of neighboring spins increases. The key insight is that fluctuations in the effective field experienced by each spin depend not just on the geometric coordination number, but on the relative strengths of different exchange pathways. For a given spin, the effective field is $\sum_{\delta}J_{\delta}\mathbf{S}_{i+\delta}$, where the sum runs over all neighbors. The relative fluctuation of this field should scale with the standard deviation divided by the mean value.
We can draw an analogy to a binary random distribution (which would correspond to spin 1/2). It is known that after $N$ attempts in a binary random distribution $x$ with probabilities $p$ and 1p1-p for the values $x=1$ and x=1x=-1, respectively, the mean value (proportional to the magnetic moment) is $\langle x\rangle=(2p-1)N$, and the standard deviation is $\sqrt{\langle x^{2}\rangle-\langle x\rangle^{2}}=2\sqrt{Np(1-p)}$. Thus, the ratio between the standard deviation and the mean value decreases as $1/\sqrt{N}$. Similarly, in our quantum Heisenberg model, one expects that as the number of neighbors increases, the specific heat is more similar to the mean-field result, in which spatial and quantum fluctuations are neglected and the effective magnetic field that each ion feels is replaced by its mean value.
Generalizing the above results to the case of a distribution with intensities $\pm|J_{\delta}|$ for $z_{\delta}$ neighbors, we obtain after some algebra (see Appendix):
$\frac{\langle x^{2}\rangle-\langle x\rangle^{2}}{\langle x\rangle^{2}}=\frac{4p(1-p)}{(2p-1)^{2}}\frac{1}{z},\quad\mathrm{where}\quad\frac{1}{z}=\frac{\sum_{\delta}z_{\delta}J_{\delta}^{2}}{\left(\sum_{\delta}z_{\delta}|J_{\delta}|\right)^{2}},$ (4)
Here, $z_{\delta}$ represents the number of neighbors of type $\delta$ (with exchange interaction $J_{\delta}$), and the summation goes over all types of neighbors. The quantity $z$ quantifies the effective number of neighbors, taking into account the strength ($|J_{\delta}|$) and number ($z_{\delta}$) of each type of interaction. Remarkably, for unfrustrated systems, the observed specific heat depends mainly on this effective number of neighbors $z$, largely independent of the specific lattice structure. $z$ quantifies the relative strength of spatial fluctuations. For larger $z$ (higher coordination number), the results approach the mean-field ones, where spatial and quantum fluctuations are suppressed. Conversely, for lower $z$ (fewer neighbors), fluctuations become more important, leading to a smaller specific heat below the critical temperature $T_{C}$ and a larger specific heat above $T_{C}$. Notably even for $z$ as large as $26$ there is a large departure of the quantum specific heat respect to the mean field result.
I.2 Axial anisotropy in 4f7 ions
Since neither the spin nor terms with $L=0$ are affected by crystal fields, one expects that the only source of splitting of the ground-state octuplet is the exchange interaction between rare-earth ions responsible for the magnetic structure of the compound. This holds as a first approximation. In fact, we will show that the specific heat of five Gd compounds can be well fitted ignoring crystal-field effects.
Nevertheless, due to spin-orbit coupling, the ground-state multiplet acquires a small admixture of the $^{6}P_{7/2}$ state with $L=1$. An explicit calculation for Gd+3 in Appendix B of Ref. 18 estimates that the excited multiplet contributes approximately $3\%$ to the ground state. This admixture introduces a component with $L=1$ in the ground state. It allows the ground state to split under the influence of crystal fields. This occurs if the symmetry at the rare-earth site is not high. Specifically, if there is only a single axis of higher order, the anisotropy term $H_{A}=C[3L_{z}^{2}-L(L+1)]$ becomes relevant.
In high-symmetry environments such as tetragonal, cubic, or octahedral point groups ($T$, $T_{d}$, $T_{h}$, $O$, $O_{h}$), the $L=1$ states remain unsplit. This result, derived from group theory, is intuitively understood. A basis for $L=1$ states can be chosen to transform like the unit vectors $\hat{x}$, $\hat{y}$, $\hat{z}$, which are equivalent in high-symmetric environments. In contrast, if the maximum order axis of the point group is 2, a term $L^{2}_{x}-L^{2}_{y}$ is also allowed. In even lower symmetries, $L_{x}L_{y}$ may arise. However, their intensity is expected to be much smaller. Thus, experimental data can be fitted without including them.
Using the Wigner-Eckart theorem, $H_{A}$ can be written in the form $H_{A}=K[3J^{2}_{z}-J(J+1)]$, where $K$ is of order $\sim 3\%C$.
II Methods
We have calculated the specific heat of a large number of magnetic structures. The approach used a two-point numerical differentiation to compute the derivative of the internal energy as a function of temperature. The internal energy was calculated using quantum Monte Carlo (QMC) simulations with the default “minbounce” algorithm (implemented in the “dirloop_sse” application in the ALPS libraries [19, 20]). QMC allow us to handle large system sizes and complex lattices while properly accounting for spatial and quantum fluctuations that mean-field approaches miss. Simulations were performed on systems of up to $8\times 30^{3}=216000$ magnetic moments. A typical simulation required approximately 10000 CPU hours per lattice on a standard workstation.
To generate the lattice configurations for the ALPS simulations, we employed the “SUNNY” package [21] in conjunction with crystallographic information files (CIF) obtained from public repositories. Specifically, CIF files for Fe in simple cubic (SC), body-centered cubic (BCC), hexagonal close-packed (HCP), and face-centered cubic (FCC) structures, as well as C in diamond, MgAgAs (for GdPbBi), Th(CrSi)2 (for GdCu2Ge2), ErNiSi3 (for GdNiSi3), Mg2MnGa3 (for Eu2Pd3Sn) and MgAl2Cu (for EuPdSn2) were retrieved from the Materials Project [22] database. The ErNi3Al9 CIF file was obtained from the Crystallography Open Database [23] using data from reference 24. Lattice parameters for GdCu2Ge2 were taken from [25], for GdPbBi from [26], for GdNiSi3 from [27], for Eu2Pd3Sn from [28], and for EuPdSn2 from [29]. In each case, the magnetic structure was constructed by assigning a spin moment $S=7/2$ to the respective atoms (C, Fe, Mg, Th, or Er). For all lattices, except HCP, we used a rectangular cuboid supercell geometry as the conventional unit cell, resulting in varying numbers of sites within the supercell, with the diamond lattice requiring the largest (8 sites). Simulations were performed on systems containing $N^{3}$ supercells with periodic boundary conditions, typically with $N=10$, 2020, and $30$. For the critical temperature, results were extrapolated to the thermodynamic limit assuming a $1/N$ dependence.
The well-known sign problem [30] associated with QMC simulations limited our investigation to non-frustrated magnetic structures. At low temperatures, the number of Monte Carlo steps required for statistically reliable results increases exponentially. Consequently, the internal energy was primarily computed in a temperature range between $0.2T_{C}$ and $2T_{C}$, where $T_{C}$ is the critical temperature of the system. In a few cases, simulations were extended to $0.1T_{C}$ at low temperatures. However, the reliability of the results below $0.2T_{C}$ is limited due to finite-size effects. We used $2\times 10^{5}$ Monte Carlo steps for thermal equilibration and $4\times 10^{6}$ steps for measurements. These parameters results in a well converged energy and a smooth specific heat. The resulting error bars (not shown in figures) are close to the line width or fluctuations seen in figure 2.
The coordination number (number of neighbors included) for each lattice is provided in parentheses after its name. For example, FCC(18) refers to a face-centered cubic lattice with 12 first neighbors and 6 second-nearest neighbors. Additional neighbor shells were included by increasing the cutoff radius $d$ for neighbor interactions.
Our theoretical results for the specific heat of all lattices used here are available in Ref. 31.
III Role of the effective number of neighbors
Figure 1: Critical temperature $T_{C}$ as a function of effective number of neighbors $z$ for several ferromagnetic structures. Finite-size results ($N=10,20,30$) and extrapolated values are shown, with the mean-field value $T_{C}^{\mathrm{MF}}$ indicated by the straight line.
In this section we report our results for the specific heat and critical temperature for several simple lattices as a function of the number of neighbors that interact with a given site. For simplicity, in all lattices considered, all sites are equivalent, and we assume all $J_{\delta}=J_{1}$, where $J_{1}$ is the nearest-neighbor (NN) interaction, up to a certain distance $d$ and 0 for longer distances. Therefore, the coordination number given by equation 4 reduces to $\Sigma_{\delta}z_{\delta}. zz$ is given by the total number of neighbors within a sphere of radius $d$. In Section V we discuss more general cases demonstrating that the results are not severely affected by this assumption.
Most of our studies assume ferromagnetic interactions ($J_{\delta}<0$). Generally, the results for the corresponding non-frustrated antiferromagnetic structures are very similar, particularly for $z>6$, except at very low temperatures, where the specific heat, dominated by magnons, scales as $T^{3/2}$ ($T^{3}$) for ferro- and antiferromagnetic structures respectively.
In Fig. 1, we show the critical temperature $T_{C}$ as a function of $z$ for several ferromagnetic structures, including finite-size results and an extrapolated value. Calculations have been done in finite systems containing $N^{3}$ conventional unit cells, with $N=10$, 20, and 30, as explained in Section II. The extrapolated value is also indicated in the figure.
As expected, $T_{C}$ increases almost linearly with $z$ as the mean-field value $T_{C}^{\mathrm{MF}}$ (the effective magnetic field due to the interactions is proportional to $zJ_{1}$, see Section IV.1). But it is lower than $T_{C}^{\mathrm{MF}}$, particularly for low $z$ due to the effect of spatial and quantum fluctuations, which is larger for low $z$.
For large enough $z$, one expects that $T_{C}$ approaches $T_{C}^{\mathrm{MF}}$. However, in finite systems with periodic boundary conditions, the coordination number decreases when the interaction range $d$ becomes comparable to the system size, because neighbors at different distances overlap. This might prevent us from reaching the thermodynamic limit with our method. Nevertheless, arguing as in the introduction that fluctuations decrease as $1/z$ , we should expect that the critical temperature approaches the MF result much faster than the results shown on Fig. 1.
Figure 2: Specific heat $C/T$ versus reduced temperature $T/T_{C}$ for several ferromagnetic structures with different effective neighbor counts $z$. Top panel ($z\leq 12$) and bottom panel ($z\geq 12$) show distinct ranges. The curves for SC, FCC, and HCP structures coincide at $z=18$, while SC, BCC, and EuPdSn2 coincide at $z=26$, explaining the single curves shown for each set. LHc(4) denotes the layered honeycomb lattice with $z=4$. The bottom inset shows finite-size effects for HCP(12) with $N=10$ (black), $N=20$ (red), and $N=30$ (green) in the range $0.5<T/T_{C}<1.5$.
In Fig. 2 we show the results for the specific heat $C$ as a function of temperature for a large number of ferromagnetic structures with different numbers of magnetic neighbors $z$ (the number of nearest neighbors up to third order is shown in Table 1). Also shown is the mean-field result which corresponds to the limit $z\rightarrow\infty$. This result shows larger $C$ for $T<T_{C}$, while $C=0$ for $T>T_{C}$. In contrast, the specific heat for a honeycomb layered system with $z=4$ (three nearest neighbors within the plane of the honeycomb layer and one above or below the layer in alternating sites, LHc(4)) shows the smallest specific heat below $T_{C}$ and the largest one above $T_{C}$. Note that the diamond structure, which is distinct from the previous one but also has $z=4$, has a very similar specific heat. The low-temperature shoulder is a Schottky-like effect due to the internal field [32, 14]. The lower panel inset shows that finite size effects are only appreciable near $T_{C}$ as exemplified in the case of the HCP(12) lattice. For this reason, we show specific heat curves from QMC corresponding to $N=20$.
Increasing $z$ from 4 to 6, the specific heat of the simple cubic structure with six NN displays the expected trend of increasing (decreasing) $C$ below (above) $T_{C}$. Increasing $z$ up to 14 using the BCC structure with 6 NN and 8 next-nearest neighbors (NNN), the same trend is observed in general. However, the specific heat of the BCC lattice with $z=8$ (first neighbors) lies very close to the cubic FCC with $z=12$ and above $z=10$. The FCC lattice with $z=12$ has a C quite close to the HCP with $z=12$ (first and second neighbors). The diamond(16) and EuPdSn2(16) lattices when 16 neighbors are considered also have very similar C. Other exceptions correspond to the FCC(18) case composed of twelve NN and six NNN that has C quite close to the previous mentioned cases of diamond(16) and EuPdSn2(16). Nevertheless, the specific heat is very similar for different structures with the same $z$. Even if the mean-field theory predicts identical behavior for all coordination numbers, without accounting for lattice geometry, this universality seen with QMC across different lattice types is remarkable.
Lattice 1st NN ($z_{1}$) 2nd NN ($z_{2}$) 3rd NN ($z_{3}$)
EuPdSn2 2 2 4
Eu2Pd3Sn3 1 2 1
GdNiSi3 2 2 2
GdCu2Ge2 4 4 8
LHc 3 1 6
Diamond 4 12 12
SC 6 12 8
BCC 8 6 12
FCC 12 6 24
HCP 12 6 2
Table 1: Number of neighbors at increasing distances for various lattice types.
Lattice $z_{1}$ $z_{2}$ $z_{3}$ $J_{1}$ $J_{2}$ $J_{3}$ $T_{C}$ (K)
LHc(4,AF) 3 1 6 1 1 0 13.2
LHc(4) 3 1 6 -1 -1 0 13.1
Diamond(4) 4 12 12 -1 0 0 13.1
SC(6) 6 12 8 -1 0 0 22.4
BCC(8,AF) 8 6 12 1 0 0 32.2
BCC(8) 8 6 12 -1 0 0 32.6
LHc(10) 3 1 6 -1 -1 -1 39.2
HCP(12) 12 6 2 -1 -1 0 49.6
FCC(12) 12 6 24 -1 0 0 49.8
BCC(14) 8 6 12 -1 -1 0 59.4
EuPdSn2(16) 2 2 4 -1 -1 -1 69.1
Diamond(16) 4 12 12 -1 -1 0 70.1
FCC(18) 12 6 24 -1 -1 0 76.9
SC(18) 6 12 8 -1 -1 0 81.4
BCC(26) 8 6 12 -1 -1 -1 122.
Table 2: Extrapolated critical temperatures as computed with QMC for Hamiltonians in the lattices considered on text. Lattices marked with * have more than third nearest-neighbors (see text). For lattices marked with we report the $N=20$ critical temperature. Uncertainty in $T_{C}$ is around 3%. correspond to a Neel temperature.
IV Effect of anisotropy
In this section, we address the role of anisotropy first in the mean-field approximation and then incorporate the effect of fluctuations.
IV.1 Mean-field Approximation
In this subsection, we discuss the effect of the term $H_{A}$ on the specific heat calculated in the mean-field approximation. In this approximation it is assumed that each site is subjected to a constant effective magnetic field resulting from the interaction with the remaining sites. Specifically, the interactions appearing in $H_{0}$ [see Eq. (1)] are approximated as
$\mathbf{S}_{i}\cdot\mathbf{S}_{j}\approx\langle\mathbf{S}_{i}\rangle\cdot\mathbf{S}_{j}+\mathbf{S}_{i}\cdot\langle\mathbf{S}_{j}\rangle-\langle\mathbf{S}_{i}\rangle\cdot\langle\mathbf{S}_{j}\rangle,$ (5)
The mean-field approximation assumes that all sites are equivalent (long-range magnetic order), which may not hold near the ordering temperature. We assume that there is no external magnetic field. For simplicity, we restrict the analysis to the easy axis (uniaxial anisotropy or anisotropy along a single axis) case $K<0$, for which the spins point in the direction of the anisotropy axis $z$. In the next subsection, both signs of $K$ and the effects of spatial and quantum fluctuations will be considered.
For an antiferromagnetic order, we rotate 180 degrees the spins pointing in the z-z direction around a perpendicular axis so that for all sites $\langle S_{z}\rangle\geq 0$. This expectation value serves as the order parameter for the magnetic phase.
The effective magnetic field $B_{\mathrm{eff}}$ is given by
$Z$ $=\sum\limits_{M=-J}^{J}e^{-\beta E_{M}}\mathrm{,~}\langle S_{z}\rangle$ $=\sum\limits_{M=-J}^{J}\frac{Me^{-\beta E_{M}}}{Z}\mathrm{,~}$
$\tilde{B}$ $=g\mu_{B}B_{\mathrm{eff}}=I\langle S_{z}\rangle,$ (6)
where $g$ is the gyromagnetic factor, $\mu_{B}$ is the Bohr magneton, $I=|\Sigma_{\delta}z_{\delta}J_{\delta}|$ is the sum of all interactions ($z_{\delta}$ is the number of neighbors at distance $\delta$ and $J_{\delta}$ its intensity), $\beta=1/(k_{B}T)$ where $T$ is the temperature and $k_{B}$ is the Boltzmann constant. The energies $E_{M}$ for each eigenvalue $M$ of the operator $S_{z}$ are given by:
$E_{M}=-\tilde{B}M+K[3M^{2}-S(S+1)].$ (7)
Eqs. (6) and (7) permit to determine $\tilde{B}$ self-consistently. In the limit $\tilde{B}\rightarrow 0$ one obtains an equation for the critical temperature
$k_{B}T_{C}^{\mathrm{MF}}=I\langle S^{2}_{z}\rangle\mathrm{,where~~}\langle S^{2}_{z}\rangle=\sum\limits_{M=-J}^{J}\frac{M^{2}e^{-\beta E_{M}}}{Z}.$ (8)
The specific heat is obtained from the numerical derivative of the entropy $S$ with respect to temperature:
$C=T\frac{dS}{dT}\mathrm{,~~~}S=\sum\limits_{M=-J}^{J}p_{M}\ln p_{M}\mathrm{,~~~}p_{M}=\frac{e^{-\beta E_{M}}}{Z}$ (9)
Figure 3: Specific heat $C/T$ versus $T/T_{C}$ for the mean-field approximation with varying axial anisotropy $K$ (expressed in units of $k_{B}T_{C}$).
In Fig. 3 we show the resulting effect of the uniaxial anisotropy on the specific heat keeping the value of $k_{B}T_{C}=1$ constant, as the unit of energy. For $K=0$, the value of the intensity for a given $T_{C}$ is $I=3k_{B}T_{C}/[S(S+1)]$, leading to $I=4/21\sim 0.190$ for $k_{B}T_{C}=1$.
For K=0.02K=-0.02, we obtain $I=0.1531$ from Eq. (8). This is because the expectation value $\langle S^{2}_{z}\rangle$ that enters Eq. (8) is larger for $K<0$ and should be compensated with a smaller $I$. In addition, since the values of high $S_{z}$ projection $|M|$ are favored, the effective magnetic field is more efficient than for $K=0$ in splitting the lowest energy levels and therefore the entropy and specific heat decrease for $T<0.4T_{C}$ with respect to the isotropic case. This is compensated by an increase in $C/T$ for $0.5T_{C}<T<T_{C}$.
For K=0.04K=-0.04, II decreases further to $I=0.1263$, and the changes in the specific heat are more marked in the same direction. For large negative $K$, $I=1/(7/2)^{2}\sim 0.0816$.
In the mean-field approximation, anisotropy shifts entropy from low to intermediate temperatures, with $T_{C}$ determined by the balance between exchange interactions and anisotropy. Above $T_{C}$ and when $K\neq 0$ the specific heat is different from zero reflecting the onsite energy fluctuations induced by the CEF term.
IV.2 Beyond Mean-field
Figure 4: Specific heat $C/T$ versus $T/T_{C}$ for the EuPdSn2(16) structure with varying axial anisotropy $K$ ($J_{\delta}=-1$). The critical temperatures for K=0.6,0.3,0,0.3,0.6K=-0.6,-0.3,0,0.3,0.6 are approximately $78,74,68.5,71.5,73$ respectively. To compare with Figure 3 the ratio $K/{k_{B}T_{c}}$ for these cases are .0077-.0077, 0.0040-0.0040,0,$0.0042$ and $0.0082$ respectively.
In this subsection, we discuss the effect of anisotropy in cases where the fluctuations of the effective magnetic field are also included. Figure 4 shows the specific heat for the EuPdSn2(16) structure (described in section VI) with varying anisotropy $K$ for moderate values ($|K|\lesssim J_{\delta}$). For $K=0$, the system is isotropic, while negative $K$ values introduce axial anisotropy, favoring alignment along the $z$-axis. $K>0$ introduces a plane anisotropy favoring alignment in the $xy$-plane. Qualitatively, the results for $T<T_{C}$ mirror those of the mean-field approximation, with anisotropy shifting entropy from the region $T<0.5T_{C}$ to $0.5T_{C}<T<T_{C}$, and the effect is more pronounced for negative $K$.
An anisotropy-induced energy gap arises from the splitting of spin states due to the axial anisotropy term $H_{A}=K[3S_{z}^{2}-S(S+1)]$. For negative $K$, this term favors alignment along the $z$-axis, creating a gap between low-energy $|S_{z}|$ states with large $|S_{z}|$ values. This restricts spin fluctuations above $T_{C}$, reducing entropy and specific heat. For positive $K$, the anisotropy favors in-plane alignment, leading to a inverse distribution of low-energy states but with similar effects on the specific heat. The QMC simulations explicitly capture these quantum and spatial fluctuations, which mean-field theory cannot account for.
The Schottky-like shoulder observed in the specific heat is attributed to the internal magnetic field which creates a low-energy gap in the spin excitation spectrum. This feature is particularly sensitive to the axial anisotropy. For positive $K$, the low-energy states are more closely spaced, leading to a smaller effect compared to negative $K$ where low-energy states are more separated. In this last case the Schottky-like shoulder is more affected by the anisotropy.
The entropy removed above $T_{C}$ by the anisotropy term is recovered between the Schottky-like shoulder and $T_{C}$.
V Comparison of Different Structures with the Same Effective Number of Neighbors
Figure 5: Specific heat $C/T$ versus $T/T_{C}$ for various lattice structures with different values of effective neighbor count $z$ (defined by Eq. 4), with $J_{1}=-1$. The critical temperatures $T_{C}$ and $z$ values are tabulated in Table 3. Top panel: Diamond lattice with varying $J_{2}$ interactions ($z=4,6,8,10,12,14,16$ from bottom to top). Inset: $z$ as a function of $J_{2}$. Middle panel: Structures with $z=8$ or $12$ (diamond, BCC and FCC). Bottom panel: Structures with $z=6$ or $18$ (diamond, SC, BCC and FCC).
Our central finding is that the specific heat of magnetic systems is primarily determined by the effective number of neighbors $z$, rather than the specific lattice structure or interaction pattern. In Section III, we analyzed the impact of varying the number of neighbors $z$ on the specific heat, while keeping $J_{\delta}=J_{1}$. Here, we extend this analysis by investigating how changes in $J_{\delta}$ affect the specific heat, while maintaining the effective number of neighbors $z$ constant [see Eq. (4)]. The parameter $z$ quantifies the relative strength of spatial and quantum fluctuations. Physically, $z$ represents the effective number of neighbors that meaningfully contribute to the exchange field experienced by each spin, accounting for both geometric coordination and relative interaction strengths.
To illustrate this principle, we use a diamond lattice with interactions $J_{1}=-1$ and varying $J_{2}$ as a guide. This allows us to explore cases where $z$ varies between $4$ (just nearest neighbors) and $16$ (nearest and next-nearest neighbors). In Figure 5(a), we show how the specific heat changes with $J_{2}$, with $z$ values ranging from $4$ to $16$. The specific heat curves closely follow the results shown in Figure 2, with even the $z=8$ and $z=12$ cases being nearly degenerate.
In the middle panel (b) and bottom panel (c) of Figure 5, we compare these diamond and non-uniform BCC and FCC results against lattices with uniform couplings (SC(6), BCC(8), FCC(12), FCC(18)). The specific heat curves for diamond and uniform coupling lattices with the same $z$ are strikingly similar. Also the diamond lattice with $J_{2}=-0.155$ ($z=8$) matches the BCC lattice with $J_{2}=-8.0$ ($z=8$), and the diamond lattice with $J_{2}=-0.333$ ($z=12$) matches the FCC lattice with $J_{2}=-4.0$ ($z=12$) . Despite the drastic changes in interaction magnitudes, the specific heat curves align closely, underscoring the dominance of $z$ over detailed interaction strengths. Even for the FCC lattice with $J_{2}=-4$, where curves do not overlap, the comparison remains good.
V.1 Getting information beyond mean field
The key insight of our work is that the effective number of neighbors $z$ (defined in Eq. 4) is the dominant parameter determining the specific heat, rather than the specific lattice structure or interaction pattern. This allows us to extract valuable information about the underlying exchange interactions that goes beyond what can be obtained from mean-field theory alone.
To illustrate this, consider the diamond lattice with varying $J_{2}$ (Fig. 5). For different values of $J_{2}$, we obtain different effective number of neighbors $z$. Crucially, when we compare diamond lattices with different $J_{2}$ values but the same $z$ to reference lattices with uniform couplings (SC, BCC, FCC), we find that the specific heat curves match almost perfectly. This demonstrates that the specific heat is insensitive to the detailed interaction pattern as long as $z$ is the same.
This observation provides a powerful tool for experimentalists. By comparing experimental specific heat data to our reference curves (available in Ref. 31), one can extract the effective number of neighbors $z$ and, from $z$, infer the relative strengths of different exchange interactions. This is a significant advance over mean-field analysis, which cannot distinguish between different interaction patterns that yield the same critical temperature.
Let’s examine how this works in practice using the examples from Fig. 5:
  1. 1.
    For the diamond lattice with $J_{2}=-0.080$ ($z=6.03$), the specific heat matches SC(6). Using Eq. 4 with $z_{1}=4$ and $z_{2}=12$ for the diamond lattice, we can solve for $J_{2}/J_{1}$:
    $1/z=\frac{z_{1}J_{1}^{2}+z_{2}J_{2}^{2}}{(z_{1}J_{1}+z_{2}J_{2})^{2}}=1/6.$
    With $J_{1}=-1$, this yields $J_{2}/J_{1}\approx 0.0787$, which is very close to the actual value of $0.080$. This demonstrates that the effective $z$ parameter can be used to extract the ratio of interaction strengths with high accuracy. Using the critical temperature as an energy scale, we can even determine the absolute values of $J_{1}$ and $J_{2}$.
  2. 2.
    Diamond lattices with $J_{2}=-0.155$ ($z=8$) and $J_{2}=-0.333$ ($z=12$) both match BCC(8) and FCC(12) as specific heat with $z=8$ and $12$ are very close. This indicates that these lattices require next-nearest-neighbor interactions for an accurate description. The analysis limits the possible parameter sets to $J_{2}/J_{1}\approx 0.155$ or $J_{2}/J_{1}\approx 0.333$, providing valuable constraints on the interaction strengths.
  3. 3.
    As in the previous case BCC($J_{2}=-0.40$) and BCC($J_{2}=-8.0$) lead to multiple possibilities: if $z=8$, then $J_{2}/J_{1}=0$ or $8$; if $z=12$, then $J_{2}/J_{1}=0.390$ or $2.276$.
  4. 4.
    For FCC($J_{2}=-4.0$), only two possibilities exist: $J_{2}/J_{1}=4$ ($z=12$) or $J_{2}/J_{1}=12.3$ ($z=8$).
  5. 5.
    For BCC($J_{1}=J_{2}=-1.0$, $J_{3}=-0.175$), the specific heat matches FCC(18) and lattices with $z=16$. This indicates that both nearest-neighbor (8) and next-nearest-neighbor (6) interactions are relevant. Considering interactions up to third order leads to a continuum of possibilities that can be further constrained by additional data.
The remarkable universality of specific heat curves across different lattice types with the same z suggests that the thermodynamics of these systems is primarily governed by the statistical properties of the exchange field distribution rather than detailed lattice geometry. As long as the exchange interactions are unfrustrated and the effective field distribution has the same variance-to-mean ratio (characterized by z), the thermodynamic behavior becomes universal.
This approach provides a systematic way to extract information about the underlying exchange interactions from specific heat measurements. It goes beyond mean-field theory by providing bounds for the number of relevant couplings and, in some cases, a limited set of options for interaction strengths. This is particularly valuable for experimentalists who often have limited information about the detailed magnetic interactions in their materials.
Lattice $J_{1}$ $J_{2}$ $J_{3}$ $z$ $T_{C}$ (K)
Diamond -1.0 -0.080 0.0 6.03 18.1
Diamond -1.0 -0.155 0.0 8 23.1
Diamond -1.0 -0.236 0.0 10 27.6
Diamond -1.0 -0.333 0.0 12 33.3
Diamond -1.0 -0.472 0.0 14 41.4
Diamond -1.0 -3.527 0.0 14 201.
BCC -1.0 -8.0 0.0 8 231.
BCC -1.0 -0.40 0.0 12 43.1
BCC -1.0 -1.0 -0.175 18 70.4
FCC -1.0 -4.0 0.0 12 151.
Table 3: Critical temperatures for Hamiltonian models with unequal interactions. All data corresponds to $N=20$.
The practical significance of this finding is substantial. By comparing experimental specific heat to our reference curves, one can extract the effective number of neighbors $z$ and, from $z$, infer the relative strengths of different exchange interactions. This information is inaccessible through mean-field analysis alone and can guide the development of more accurate theoretical models. For example, if a specific heat of a SC compound matches our reference curve for $z=8$ (and so it also matches $z=12$), we can immediately infer that the system likely has a combination of nearest-neighbor and next-nearest-neighbor interactions with a ratio $J_{2}/J_{1}\approx 0.08$ ($z=8$) or $J_{2}/J_{1}\approx 0.25$ ($z=12$).
This framework represents a significant advance in our ability to interpret magnetic specific heat data for 4f7 systems, as demonstrated in Section VI where we apply these principles to fit experimental data for six different compounds.
VI Fits of experimental data
In this section, we present fits of the specific heat of six compounds: GdNi3Ga9, GdCu2Ge2, GdPbBi, GdNiSi3, Eu2Pd3Sn3, and EuPdSn2 (see Figure 6). The fitting procedure uses the experimentally determined critical temperature $T_{N}$ as a fixed input, with the effective number of neighbors $z$ and axial anisotropy $K$ as free parameters. For Gd compounds, $K=0$ was used as the anisotropy term was found to be negligible. The assumption of uniform exchange interactions ($J_{\delta}=J$) is justified by our theoretical results showing that different $J_{\delta}$ values with the same effective number of neighbors $z$ yield nearly identical specific heat curves (see Section V).
The quality of the fits is excellent. For comparison, a mean-field approach would fail to reproduce the specific heat curves, particularly the shape of the peak and the high-temperature tail (see MF specific heat in Figures 2 and 3).
The parameters of the fits for the different compounds are listed in Table 4. The table also shows $z_{0}$. It represents the minimum number of nearest neighbors required to maintain 3D connectivity in the reported crystal structure. The effective number of neighbors $z$ accounts for both geometric coordination and relative interaction strengths, which can differ from the nominal $z_{0}$ due to longer-range interactions.
Compound $T_{N}$ (exp,[K]) $_{0} zz$(fit) |J||J|[K] KK[K]
GdNi3Ga9 20 4 4 1.5 0
GdCu2Ge2 12 12 12 0.20 0
GdPbBi 13 12 12 0.26 0
GdNiSi3 22 17 33 0.15 0
Eu2Pd3Sn 47 4 14 0.71 -0.28
EuPdSn2 12 8 22 0.11 -0.09
Table 4: Parameters used in the fits: $T_{N}$ is the Neel temperature, $z_{0}$ is the number of nearest neighbors in the reported structure (compatible with 3D connectivity), $z$ (fit) is the effective number of neighbors, $J$ is the interaction strength, and $K$ is the anisotropy. This value correspond to $J_{1}=J_{2}$ and $J_{3}=0.27J_{1}$, see main text.
GdNi3Ga9 [33, 34] exhibits a layered honeycomb structure with $z_{0}=4$ (three intralayer and one interlayer nearest neighbors), consistent with $z=4$ found in the fit. The fit (Figure 6a) reproduces the experimental specific heat curve with excellent accuracy, with the peak at $T_{C}=20$ K matching the experimental value. The small $z$ value indicates significant spatial and quantum fluctuations, which is consistent with the observed broadening of the specific heat peak. As LHc(4,AF) lattice has $T_{N}/{|J_{1}|}\sim 13$ while experimentally $T_{N}=20K$ we can deduce $J\sim 20$$13\sim 1.5K$ corresponding to an antiferromagnetic interaction.
For the rest of the structures, ferromagnetic interactions are assumed. As stated in Section III, as long as there is no significant frustration, the difference between ferromagnetic and antiferromagnetic interactions for $z>6$ is only noticeable at very low temperatures, well below the temperature range of interest in this study. In figure 5 we show the specific heat of a BCC lattice with AF interactions that compares well with the FM case. The key features of the specific heat curves, such as the shape of the peak and the high-temperature tail, are primarily determined by the effective number of neighbors and the single-ion anisotropy. These features are largely insensitive to the sign of the exchange interaction in the absence of strong frustration. For this reason we compare the experimental Neel temperature ($T_{N}$) with the theoretical Curie temperature ($T_{C}$).
GdCu2Ge2 has a tetragonal structure[35]. The Gd sublattice adopts a body-centered tetragonal structure. As this lattice is quite elongated ($c/a\gg 1$) if only the first four nearest-neighbor interactions were considered, the system would become two-dimensional with a broadened transition near $T_{N}$. The four next-nearest neighbors still lay in the same plane as the NN. To recover a tridimensional connectivity compatible with the experiments at least the eight following neighbors must be considered. This leaves as a minimum option to consider the first and third nearest neighbors ($z_{0}=12$) interactions. As explained before, this should be similar to consider other lattice with $z=12$ as the FCC lattice. In panel b) we show that indeed the specific heat can be described by any of these lattices. If, to simplify, we consider $J_{1}$ ($z_{1}=4$) equal to $J_{2}$ ($z_{2}=4$) we get the equation
$1/z=\frac{(z_{1}+z_{2})J_{1}^{2}+z_{3}J_{3}^{2}}{((z_{1}+z_{2})J_{1}+z_{3}J_{3})^{2}}.$
This assumption leads to three possibilities:
  • $J_{3}=0$ ($z=8$) which we can disregard has it implies a two dimensional behavior
  • $J_{3}/J_{1}=3.7$ ($z=12$) which we can disregard has it seems unphysical to have a larger strength associated to a more distant bond.
  • $J_{3}/J_{1}=0.27$ ($z=12$) appears to be the only sensitive solution.
Without the actual simulation for this combination of couplings, we can only estimate the coupling’s strength. To get the strength of the coupling we use the MF result for the critical temperature (see section IV.1)
$\frac{k_{B}T_{C}}{\frac{S(S+1)}{3}}=(z_{1}+z_{2})|J_{1}|+z_{3}|J_{3}|=(8+8\times 0.27)|J_{1}|$
from which we obtain $|J_{1}|=|J_{2}|\sim 0.20$K and $|J_{3}|\sim 0.05$K. From this results we can not determine the signs of the coupling. But this antiferromagnetic compound at least one of these coupling should be antiferromagnetic.
GdPdBi [36] is a half-Heusler structure. Gd sublattice in this compound forms a FCC structure. We obtain a very good fit with the LHc(10) lattice. Using $z=10$ leads to $J2/J_{1}\sim 6.1$ which seems unphysical. As FCC(12) ($T_{C}/{|J_{1}|}\sim 50$) also accounts for the experimental data we use that lattice to estimate the interaction strength. As in the previous compound we suppose that a NN interaction alone scenario is closer to the compound. Experimentally $T_{N}=13$K so we deduce $|J|\sim 13$K/50=0.26/50=0.26K. For this antiferromagnetic compound an antiferromagnetic coupling is expected.
The crystal structure of GdNiSi3 is an orthorhombic system. Gd atoms forms a bilayered arrangement of Gd. Only interaction at the 7th nearest neighbor connects bilayers to provide a tridimensional connectivity ($z_{0}=17$). A good description of the experimental data is obtain considering up to 10th nearest-neighbor ($z=25$) while the best fit is seen considering up to 13th NN ($z=33$). This suggests that long range interactions, well beyond first neighbors, are relevant in this compound. Given the complexity of these interactions we only consider the uniform $J_{\delta}$ case. The GdNiSi3(33) lattice has $T_{C}/{|J_{1}|}\sim 145$ while experimentally $T_{N}=22$K so we can deduce $|J|\sim 22$K/145=0.15/145=0.15K. The large amount of couplings needed (up to 13 nearest-neighbors) certainly come from a distribution of interactions. This distribution could be determined with first principle calculations that considers at least this coordination.
This completes the Gd compounds that we have considered. In a previous work [32], the specific heat of GdCoIn5 has been fitted with essentially the same procedure. The compound has a tetragonal structure. The authors have extrapolated to the thermodynamic limit the specific heat that corresponds to a simple cubic lattice $z=6$ nearest-neighbor with antiferromagnetic interactions, providing an excellent fit of the data.
Now turning to Eu compounds, the importance of $K$ for the Eu compounds may be attributed to the $\sim 25\%$ larger ionic radius of Eu2+ compared to Gd3+ [37], which makes the former more susceptible to crystal-field effects.
The Eu sublattice of Eu2Pd3Sn can be though as a stacking of corrugated hexagonal layers in the ab plane. The smaller coordination number to have a tridimensional connectivity is $z_{0}=4$. This involves up to third order nearest-neighbor interactions. To properly compare the specific heat of Eu2Pd3Sn [38] with a Hamiltonian model we have to consider up to 6th nearest-neighbor in this lattice leading to a coordination $z=14$. The point group at the Eu sites is $C_{1h}$, which allows for the inclusion of the anisotropy term $H_{A}$, as discussed in section I.2. So an anisotropy term with $K/|J_{1}|=-0.4$ was included. With this anisotropy the Eu2Pd3Sn(14) lattice has $T_{C}/{|J_{0}|}\sim 66$. Experimentally $T_{N}=47$K so we can deduce, assuming uniform couplings, $|J|\sim 47$K/66=0.71/66=0.71K.
Finally, the Eu sublattice of EuPdSn2 can be though as a stacking of rectangular layers in the ac plane. The smaller coordination number to have a tridimensional connectivity is $z_{0}=8$. This involves up to 3 nearest-neighbor interactions. To properly compare the specific heat of EuPdSn2 [29] with a Hamiltonian model we have to consider up to 7th nearest-neighbor in this lattice leading to a coordination $z=22$. The point group at the Eu sites is $C_{2v}$, which also allows for the inclusion of the term $H_{A}$. So an anisotropy term with $K/|J_{0}|=-0.9$ was included. With this anisotropy the EuPdSn2 lattice has $T_{C}/{|J_{0}|}\sim 112$. Experimentally $T_{N}=12$K from which we deduce $|J|\sim 12$K/112=0.11/112=0.11K. The discrepancies near $T_{N}$ likely originate from factors not explicitly included in our model, such as more complex magnetic ordering (e.g., a collinear amplitude-modulated structure) resulting from modulated interactions and crystal-field effects, as discussed in Ref. 14.
Since all the compounds contain transition-metal ions, which may be magnetic, one might question their influence on the specific heat. However, our analysis shows that the entropy obtained by integrating $C/T$ is close to $\ln(8)$ per rare-earth ion (consistent with $S=7/2$), indicating that any effects from transition metals are negligible.
The success of these fits suggests that the effective number of neighbors $z$ is the dominant parameter determining the specific heat of 4f7 systems, with axial anisotropy $K$ providing additional refinement for Eu compounds. This approach significantly outperforms mean-field theory, which would fail to reproduce the observed specific heat curves (see Figure 2).
Figure 6: Experimental specific heat $C/T$ versus $T/T_{N}$ for six 4f7 compounds (dots) and our theoretical fits (lines). $K=0$ for all Gd-based compounds (GdNi3Ga9, GdCu2Ge2, GdPbBi, GdNiSi3). Experimental data sources: Refs. 33, 34, 36, 35, 27, 29, 38.
VII Summary
We present a unified framework for understanding the specific heat of 4f7 magnetic compounds (Gd3+ and Eu2+), characterized by a ground state with $J=7/2$ and $L=0$. Through extensive quantum Monte Carlo simulations, we demonstrate that the specific heat is primarily governed by two parameters: the effective number of neighbors $z$ and axial anisotropy $K$.
The effective number of neighbors $z$ quantifies the strength of spatial and quantum fluctuations, with larger $z$ values corresponding to reduced fluctuations and specific heat curves closer to mean-field predictions. Crucially, we find that the specific heat is remarkably insensitive to the detailed lattice structure for a given $z$, enabling a universal description of diverse magnetic systems.
The origin of the anisotropy $K$ lies in a small admixture of $^{6}P_{7/2}$ in the ground-state multiplet. This admixture introduces a component with $L=1$, rendering the system sensitive to axial crystal fields.
We successfully fit the specific heat of six experimental compounds (four Gd-based and two Eu-based) using only $z$ and $K$ as free parameters, with excellent agreement between theory and experiment. Notably, axial anisotropy is essential for Eu compounds but can be neglected for Gd compounds, reflecting the larger ionic radius of Eu2+ and its greater susceptibility to crystal-field effects. This work provides a simple yet powerful tool for interpreting specific heat data in 4f7 systems, significantly outperforming mean-field approaches and enabling the extraction of (effective) fundamental magnetic interaction parameters from experimental measurements. While our two-parameter model works well for the compounds studied, it may require additional parameters for systems with strong frustration or complex magnetic order.
Experimental researchers studying new 4f7 compounds can now follow this straightforward procedure: (1) Measure the specific heat and determine $T_{N}$; (2) Compare the $C/T$ vs $T/T_{N}$ curve to our reference data in Ref. [31]; (3) Identify the $z$ value that provides the best match; (4) Use Equation 4 and known crystal structure to constrain possible exchange interaction patterns. This approach provides physically meaningful constraints on magnetic interactions. These results could also guide the development of more accurate theoretical models for compounds.
Acknowledgements.
The authors are grateful to I. Čurlík, L. de Sousa Silva, J.G.S. Duque, P.G. Pagliuso, and M. Avila for allowing us access to the original experimental data. A.A.A. acknowledges financial support provided by PICT 2020A 03661 of the Agencia I+D+i, Argentina. D.J.G. is supported by PIP 11220200101796CO of CONICET, Argentina. Computational resources were provided by the HPC cluster of the Physics Department at Centro Atómico Bariloche (CNEA), which is part of Argentina’s National High-Performance Computing System (SNCAD).
VIII Appendix
Here we calculate the mean value and fluctuations of a random variable $x$ with different intensities  [^1] The same result can be obtained considering each bond as an independent variable.. For simplicity we consider only two intensities $J_{1}$ and $J_{2}$. Extension to the general case is straightforward. We consider $z_{1}$ attempts in which the probability of the result $J_{1}$ is $p$ and that of $J_{1}$ is q=1pq=1-p. In addition, there are $z_{2}$ attempts with probability $r$ of finding the result $J_{2}$ and s=1rs=1-r for the result $J_{2}$. At the end we will take $r=p$. For non-frustrated systems, both $J_{\delta}$ can be chosen positive. Clearly, the expectation value of $x$ is
$\langle x\rangle=z_{1}J_{1}(p-q)+z_{2}J_{2}(r-s).$ (10)
On the other hand, considering the probability of each individual event, we have:
$\langle x^{2}\rangle$ $=$ $\sum\limits_{n=0}^{z_{1}}\sum\limits_{m=0}^{z_{2}}\binom{z_{1}}{n}\binom{z_{2}}{m}p^{n}q^{z_{1}-n}r^{m}s^{z_{2}-m}$ (11)
$\times\left[nJ_{1}-(z_{1}-n)J_{1}+mJ_{2}-(z_{2}-m)J_{2}\right]^{2}.$
Using the operator
$O=\left[J_{1}\left(p\frac{\partial}{\partial p}-q\frac{\partial}{\partial q}\right)+J_{2}\left(r\frac{\partial}{\partial r}-s\frac{\partial}{\partial s}\right)\right],$
$\mathrm{and}(p+q)^{z_{1}}=\sum\limits_{n=0}^{z_{1}}\binom{z_{1}}{n}p^{n}q^{z_{1}-n},$ (12)
and a similar expression for $(r+s)^{z_{2}}$, where $p$ and $q$ as well as $r$ and $s$ are taken as independent variables in the derivatives, one realizes that the expectation value Eq. (11) can be written as
$\langle x^{2}\rangle=O^{2}(p+q)^{z_{1}}(r+s)^{z_{2}}.$ (13)
Performing the calculation, we find after some algebra
$\langle x^{2}\rangle$ $=$ $J_{1}^{2}z_{1}\left[1+(z_{1}-1)(p-q)^{2}\right]+J_{2}^{2}z_{2}\left[1+(z_{2}-1)(r-s)^{2}\right]$ (14)
$+2J_{1}J_{2}z_{1}z_{2}(p-q)(r-s)$
Using Eqs. (10) and (14), setting $r=p$ and replacing q=s=1pq=s=1-p, one finally obtains
$\frac{\langle x^{2}\rangle-\langle x\rangle^{2}}{\langle x\rangle^{2}}=\frac{4p(1-p)}{(2p-1)^{2}}\frac{z_{1}J_{1}^{2}+z_{2}J_{2}^{2}}{\left(z_{1}J_{1}+z_{2}J_{2}\right)^{2}}.$ (15)
Naturally, if $\langle x\rangle=0$ as it is the case for temperatures above the critical temperature $T_{C}$, Eq. (15) gives a divergent result. In this case a meaningful comparison should involve the fluctuations above $T_{C}$ and the value of $\langle x\rangle=0$ at some point below $T_{C}$.